### v1

@@ Abstract
The welfare of a dynasty that consists of identical Bernoullian utilities subject to Cobb-Douglas production with constant labor is maximized. We blend two discounting schemes to avoid the objective function from exploding to infinity: one is the compound discounting of future utilities and the other is the binary discounting which undiscounts utilities before the planning horizon but fully discounts those that lie beyond. After presenting a general solution to the optimal schedule of consumption, we study the planning solution analytically and also numerically, to discover that the optimal planning horizon is not necessarily an infinity, even under the case of zero compound discounting. Furthermore, we study the inequality of consumptions among generations with respect to the compound discount rates and planning horizons.
@@
Dynamic Programming Sustainability Optimal Population Repugnant Conclusion Social Discounting Intergenerational Equity
@@ 1 Introduction
This study is concerned with population axiology from a dynamical perspective. Greaves (2017) delineates population axiology as the study of the conditions under which one state of affairs is better than another, when the states of affairs in question may differ over the numbers and the identities of the persons. There are several approaches to population axiology; one most commonly employed in economics is the total view (utilitarianism) that evaluates populations according to the total amount of well-being that they contain. Among the variety of principles of aggregating potential well-beings, a critical-level utilitarian (e.g., Blackorby et al, 2005) particularly seeks for a state $x$ to maximize the population value $\mathcal{\mathcal{V}}=\sum_{i\in n(x)}\left({\Upsilon}_{i}(x)-\kappa\right)$, where $n(x)$ is the set of individuals, ${\Upsilon}_{i}(x)$ is the lifetime utility of an individual $i$, and $\kappa\geq 0$ is the critical level.
[…]
Under classical utilitarianism ($\kappa=0$), $s=\bar{c}>0$ is called the well-being subsistence (Dasgupta, 2019) or consumption for neutrality, since ${\Upsilon}(s)=0$. For any consumption $c_{i}<s$, utility level of an individual $i$ is negative, meaning that $i$’s life is not worth living. Hence, it must be that $c_{i}>s$ so the lifetime utility of all individuals are positive. Note however that under critical-level utilitarianism, an individual $i$’s life contributes positively to the value of a population if $\bar{c}<c_{i}$, while negatively if $s<c_{i}<\bar{c}$, even the person leads a life worth living. For analytical purposes, we employ ${\Upsilon}(c_{i})=\log c_{i}+\kappa$ with a $\kappa$ that the well-being subsistence $s=e^{-\kappa}$ is sufficiently close to zero. The population value of critical-level utilitarianism will then be $\mathcal{\mathcal{V}}=\sum_{i=1}^{n}\log c_{i}$ where $c_{i}\in[e^{-\kappa},\infty)$ and $\bar{c}=1$, which we hereafter seek to maximize, with respect to $c_{i}$ and $n$.
@@
As we recall our primary question, let us introduce descrete time in population ethics while assuming, for sake of simplicity, that periodical number of people is fixed to one unit; that is, all (representative) individuals are identified by the period $t$ they live in. [^2] We may therefore call $t$ a generation. The population ethics problem then becomes a generation ethics problem. The critical-level utilitarian population value function now becomes $\mathcal{\mathcal{V}}=\sum_{t=0}^{n}\log c_{t}$. We note that this type of value function (based on Bernoullian utility) is widely employed in the previous Ramsey-Cass-Koopmans type of dynastic optimization models (e.g., Nordhaus, 1992; Stern, 2007), except that the future generations are discounted according to the periodical deviation from the present, and typically, the number of potential generations $n$ is left unquestioned (and hence, set to infinity). In contrast, potential generations will be accounted equally whilst the number of generations will be subject to maximization, in this study.
@@
The remainder of the paper proceeds as follows. In the following section, we first specify the model of a production economy based on Cobb-Douglas technology from a dynamical perspective. We then specify our primary question in the form of a finite horizon dynamical programming problem where the objective is to maximize the sum of utilities given the abovementioned model of production, and derive a general solution function. In Section 3 we introduce two broad settings of parameters to simplify the general solution function. The first one is the AK setting where the output elasticity of capital is set to unity. The second one is the ZD (zero discounting) setting where the discount factor is set to unity. We also study the peculiarity of the two models in regard to intergenerational equity and optimal initial action of consumption. Section 4 concludes the paper.
[…]
Figure 1: Contribution to the population value (vertical axis) under critical-level utilitarianism. Left: The critical-level consumption $\bar{c}$ contributes none to the population value, i.e., ${\Upsilon}(\bar{c})-\kappa=0$. Consumption level $c^{*}$ maximizes the critical-level utilitarian population value. Right: Well-being subsistence $s$ bears zero lifetime utility (neutrality), i.e., ${\Upsilon}(s)=0$.
2 The model
2.1 The problem of a dynasity
@@
Let us begin by postulating a two-factor Cobb-Douglas production function of constant returns to scale, as follows:
$Y_{t}={A}(K_{t})^{\theta}(L_{t})^{1-\theta}=\frac{{B}}{\theta^{\theta}(1-\theta)^{1-\theta}}(K_{t})^{\theta}(L_{t})^{1-\theta}$ (1)
@@
where $Y_{t}$ denotes total output of the economy, $K_{t}$ denotes capital, and $L_{t}$ denotes labor, all effective in period $t$. As for the parameters concerned, $\theta<1$ denotes the output elasticity of capital, ${A}>0$ denotes the level of technology termed productivity, and ${B}={A}\Theta$, where $\Theta=\theta^{\theta}(1-\theta)^{1-\theta}$, denotes the cost function-based productivity, which we otherwise call the dual productivity.333 See Appendix 1 for more details.
@@
Below describes the breakdown of total output into investment, depreciation of capital, and final consumption:
@@
Here, $\delta\leq 1$ denotes depreciation factor of capital, and $C_{t}$ denotes consumption in period $t$. Combining equations (1) and (2), and dividing both sides by $L_{t}=L$ which we assume to be constant over time, leads to the following intertemporal dynamics of capital intensity:
@@
where $k_{t}=K_{t}/L$ and $k_{t+1}=K_{t+1}/L$ denote capital intensities in period $t$ and $t+1$, respectively. Also, $y_{t}=Y_{t}/L$ and $c_{t}=C_{t}/L$ denote per caput output and consumption, respectively, in period $t$.
@@ 2.2 Finite horizon dynamic programming
Our primary problem here is to solve for an optimal consumption path given any planning horizon. The Bellman equation of the problem is of the following:555 Note that the objective function of the primary problem (2.1) is given at $t=0$, i.e., $\mathcal{V}[n]=\mathcal{V}_{0}[k_{0};n]$.
@@
The optimal trajectory of the state variable $k_{t}$ is given, by Lemma 2 which we append to Appendix 2 with proofs, as follows:
@@
The following optimal trajectory of consumption is obtained by (5), and (21) which must be true by the proof of Lemma 1.
@@
With formula (7) the population value function is given as follows:
@@
We hereafter seek to maximize this function with respect to the planning horizon $n$.
Our approach to the analysis of population value (8) will be numerical rather than analytical, because the derivative for $\mathcal{V}[n]$ with respect to $n$ seems uninformative. As it turns out in the following section, the optimal tragectory of contributions $\log c_{t}^{*}[n]$ and the population value $\mathcal{V}[n]$, for any given planning horizon $n$, becomes manageable under $\theta=1$ (known as the AK setting) and $\beta<1$. We also find that the trajectory of contributions $\log c_{t}^{*}[\infty]$ and the population value $\mathcal{V}[\infty]$ for an infinite planning horizon, is evaluable under $\beta\theta<1$. We therefore base our study of zero discounting ($\beta=1$) on this setting (i.e., $\beta\theta<1$, which indicates $\theta<1$, a Cobb-Douglas model). In the following section, we delve into the abovementioned two broad settings of parameters, namely, AK production with future discounting ($\theta=1$ and $\beta<1$) which we term as AK settings, and Cobb-Douglas production without future discounting ($\theta<1$ and $\beta=1$) which we term as ZD settings.
Table 1 summarizes the parameter settings chosen for numerical examinations. Note that cases I–IV correspond to AK settings whose solution paths are characterized by the sign of $\log({A}\beta)$, whereas cases V–VII correspond to ZD settings whose solution paths are characterized by the sign of $\log(A\Theta)$. The final case VIII corresponds to the parameter settings that exemplifies $\log({A}\beta)=\log(A\Theta)=0$.
@@
case ${A}$ $\beta$ $\theta$ $\log({A}\beta)$ $\log({A\Theta})$ $n^{*}$ $\mathcal{V}[n^{*}]$ $\mathcal{V}[\infty]$
I $1.012$ $0.992$ $1$ $+$ $\infty$ $84.675$
II $1.01$ $0.992$ $1$ $+$ 95 $59.965$
III *1 $0.992$ $1$ 00 73 $55.627$
IV $1.005$ $0.992$ $1$- 58 $50.979$
@@
    For all cases I–VIII, $k_{0}=150$.
@@
    ${A}=1/\beta\approx 1.008065$.
@@
    $\theta\approx 0.955392$ where 1.2($1{1.2}(1-\theta)^{1-\theta}\theta^{\theta}=$ 1.
@@ 3.1 AK setting
AK model is one of the simplest yet fundamental models of endogenous growth, which was formally developed by Frankel (1962). Here, we take this production model to study the population value function that discounts future generations to see whether if $n\to\infty$ is an optimal policy. The optimal consumption path (7) for AK setting with future discounting ($\theta=1$, $\beta<1$) may be specified as follows:
@@
For sake of the analysis let us take the derivative with respect to ${n}$.
@@
where $\lambda$ and $\gamma$, that are both independent of $n$, are given as follows:
@@
In order to study the derivative sign of (11), we consider the following two functions whose equality convey the first order conditions, i.e., $f[n]=g[n]$.
@@
Before we proceed, let us examine the possible range of the parameter $A$. Described below is the evaluation of the marginal product of capital (MPK) of our production function (1) net of depreciation, which should coincide with the real interest rate $\rho>0$, viz.,
@@
Then, we know that AK models with complete capital depreciation $\theta=\delta=1$ amounts to ${A}=1+\rho$, and therefore ${A}>1$ is a relevant assumption in this setting.666 Note also that $A\beta=(\delta+\rho)\beta$ indicates the discrepancy between the gross rate of interest and the rate of generational preference of the population.
Figure 2 (left) depicts functions $f[n]$ and $g[n]$ under different parameters for cases I–IV. Clearly, $f[n]$ is a linear function whose slope is $\log({A}\beta)$, and $g[n]$ is monotonously increasing and approaching zero, i.e., $g[\infty]=0$. Here, we fix the discount factor $\beta$, and differentiate the productivity ${A}$ in four different levels. As long as ${A}\leq 1/\beta$, so that $\log({A}\beta)\leq 0$, i.e., the slope is zero or negative, $f[n]$ will intersect with $g[n]$ at a single point and the population value function $\mathcal{V}[n]$ has a single peak. If ${A}>1/\beta$ so that $f[n]$ has a positive slope, then $f[n]$ and $g[n]$ could either intersect with two points where the population value $\mathcal{V}[n]$ would rise, fall and then rise again, or never intersect so that the population value would be rising indefinitely, with respect to the planning horizon $n$. Figure 2 (right) depicts the population value functions for cases I–IV.
Figure 2: Left: Solid plots show $g[n]$ defined in (12) for cases I–IV with obvious correspondences, i.e., the steepest one corresponds to case I, whereas case IV corresponds to the one with a negative slope. Dashed plot shows $f[n]$ also defined in (12). Right: The plots show population value paths $\mathcal{V}[n]$ for cases I–IV with obvious correspondences.
To visualize the optimal trajectory of the variables in various situations, we specify them here in the form of functionals. By virtue of (9), the optimal trajectory of undiscounted contribution to the population value becomes linear with respect to $t$, as follows:
@@
By virtue of (21), the optimal trajectory of capital intensity becomes as follows:
@@
Figure 3 displays the population value function $\mathcal{V}[n]$, specified as (10), for cases I, III, and IV, on the top. [^7] These figures are reiterations from Figure 2 (right). Note that, as long as $\beta<1$, the population value function will always converge to some finite value, regardless of the parameter settings besides $\beta$. As we let $n\to\infty$ in (10), we have what follows:
@@
By comparing the population values at the numerical solution of the first order condition $f[n]=g[n]$ in reference to (12), with $\mathcal{V}[\infty]$ in reference to (16), we know the optimal planning horizon $n^{*}$, which we display in Table 1 for cases I–IV. At the middle row of Figure 3, optimal trajectories of undiscounted contribution to the population value, according to (14), with respect to planning horizons $n=200,400,600$, are displayed for sample I, III, and IV from left to right. Similarly at the bottom row of Figure 3, optimal trajectories of capital intensity, according to (15), with respect to planning horizons $n=200,400,600$, are displayed for cases I, III, and IV from left to right.
Figure 3: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for AK model cases I (left), III (center), and IV (right).

### v3

@@ Abstract
This paper develops a dynamic optimization framework for evaluating the optimal longevity of a dynasty under a critical-level utilitarian social welfare criterion. Unlike standard models that assume an infinite planning horizon and apply discounting to future generations, we endogenize the planning horizon itself and examine the welfare-maximizing number of generations in a roundabout production economy with capital accumulation. The model incorporates a well-being threshold that must be met for generational existence to be considered ethically justified, providing an alternative to traditional utilitarian aggregation. Using finite-horizon dynamic programming, we derive closed-form consumption trajectories and analyze how technological parameters and intertemporal preferences affect the optimal generational span. Our results show that the optimal horizon need not be infinite—even under zero discounting—and that intergenerational equity may require limiting dynasty length.
@@
Critical-level Utilitarianism Dynamic Programming Optimal Population Intergenerational Equity
@@ 1 Introduction
Economic models of intertemporal planning often assume an infinite horizon, implicitly treating dynasties or economies as perpetually forward-looking entities. However, such assumptions can obscure important trade-offs between current and future welfare, particularly when the viability of future generations depends on limited resources and capital dynamics. In this paper, we examine a social planner’s problem where the planning horizon—the number of generations considered—is endogenously chosen. By adopting a critical-level utilitarian welfare function, we evaluate the contribution of each generation relative to a minimum well-being threshold. This framework allows us to study how productivity, capital accumulation, and time preference interact to determine the ethically and economically optimal duration of a dynasty.
We begin by positing a (classical) unitilitarian social planner whose objective is to indentify a state $x$ that maximizes the sum of the utilities $\mathcal{V}=\sum_{i\in n(x)}\Upsilon_{i}(x)$, which we refer to as the population value, where $n(x)$ denotes the set of individuals and $\Upsilon_{i}(x)$ denotes the lifetime utility (i.e., the level of well-being) of an individual $i$. For simplicity, let us assume that all individuals are identical and consume an equal share, denoted by $c$, from a fixed amount of resource $R>0$. A utilitarian axiology examines the problem of finding the population size $n=R/c$ that maximizes the population value $\mathcal{V}=\sum_{i=1}^{n}\Upsilon(c_{i})=n\Upsilon(c=R/n)$. The utility function $\Upsilon(c)$ is assumed to be monotonically increasing and concave. In addition, with respect to the population value $\mathcal{V}=n\Upsilon(c)$, zero utility ($\Upsilon=0$) is equivalent to the absence of life ($n=0$).
The consumption level $c=\nu$, where $\Upsilon(\nu)=0$, be referred to as the well-being subsistence level, following Dasgupta (2019). Owing to the monotonicity of the utility function, $\Upsilon(c)>0$ must hold as long as $c>\nu$ (see Figure 1 (left)). It then follows that the domain allowed for $c$ is $(\nu,\infty)$ since $\Upsilon(c)$ represents the level of attainable lifetime utility. Figure 1 (left) depicts the optimal solution $c=\omega$, which satisfies the first-order condition for maximizing the population value, i.e., $\Upsilon^{\prime}(\omega)=\Upsilon(\omega)/\omega$. Owing to the concavity of the utility function, if the well-being subsistence approaches zero ($\nu\to 0$), as Córdoba (2023) refers to the homothetic case, the optimal consumption also approaches zero ($\omega\to\nu$), while the optimal size of the population approaches infinity ($n^{*}=R/\omega\to\infty$), leading to a conclusion that is considered repugnant (Parfit, 1986).
Meanwhile, Blackorby et al (2005) introduced a population axiology known as critical-level utilitarianism. Within this framework, a social planner aims to determine a set of consumption schedules $c_{i}$ that maximizes the population value $\mathcal{V}=\sum_{i=1}^{n}\left(\Upsilon(c_{i})-\alpha\right)$, where an individual’s contribution to the population value is evaluated relative to the critical utility level $\alpha$ (see Figure 1 (right)). That is, an individual $i$’s life contributes positively to the population value if $c_{i}>\kappa=\Upsilon^{-1}(\alpha)$ but negatively otherwise. Note that setting $\alpha>0$ prevents the repugnant conclusion even when the underlying utility function is homothetic ($\nu=0$). For convenience, we employ a utility function of the form $\Upsilon(c_{i})=\log c_{i}+\alpha$ for analyses so that the value function for the critical-level utilitarian population value in our maximization study becomes $\mathcal{V}=\sum_{i=1}^{n}\left(\Upsilon(c_{i})-\alpha\right)=\sum_{i=1}^{n}\log c_{i}$.
@@
As we recall our primary idea of incorporating roundabout production into the optimal population framework, let us introduce (discrete) time and assume that the number of individuals per period is fixed to one unit; that is, all (representative) individuals are identified by the period $t$ during which they live. [^1] We may therefore refer to $t$ as a generation. The population value function then becomes $\mathcal{\mathcal{V}}=\sum_{t=0}^{n}\log c_{t}$. Note that this type of value function was widely employed in the previous Ramsey–Cass–Koopmans type of dynastic optimization model (including Nordhaus, 1992; Stern, 2007), except that the future generations were discounted according to the temporal deviation from the present, and typically, the number of potential generations $n$ remains unquestioned (and set to infinity). In contrast, in this study, potential generations are accounted for equally, and the number of generations is optimized.
@@
The remainder of the paper proceeds as follows. In the following section, we first describe the model of roundabout production on the basis of Cobb–Douglas technology. We then define our main problem in the form of finite horizon dynamic programming, where the objective function comprises a series of Bernoullian contributions, and analytically derive a general solution. In Section 3, we introduce two broad settings of parameters to simplify the general solution and allow us to study the optimal trajectory of the variables in a numerical fashion. The first setting is the AK setting, where the output elasticity of capital is set to unity. The second is the zero discounting (ZD) setting, where the discount factor is set to unity. We also study the peculiarity of the two models with respect to intergenerational equity and the optimal initial action of consumption. Section 4 concludes the paper.
Figure 1: Left: Utility function of an individual. The consumption level for neutrality (or well-being subsistence) is denoted by $\nu$. Marginal utility and average utility coincide at $c=\omega$. Right: Contribution to the critical-level utilitarian population value, where $\alpha$ denotes the critical level. At $c=\kappa$, the utility reaches the critical level. In all cases, the domain of permissible consumption levels $c$ for existence is $c>\nu$.
2 Model
2.1 Roundabout production
@@
Let us begin by postulating a two-factor Cobb–Douglas production function characterized by constant returns to scale, as follows:
$Y_{t}={A}(K_{t})^{\theta}(L_{t})^{1-\theta}$ (1)
@@
where $Y_{t}$ denotes the economy’s output, $K_{t}$ denotes capital, and $L_{t}$ denotes labor, all of which are effective during period $t$. For the relevant parameters, $10<\theta<$ 1 denotes the output elasticity of capital. The level of technology is indicated by productivity, denoted by $A$, which is assumed to be fixed throughout the considered time span. Note that ${B}\equiv{A}\theta^{\theta}(1-\theta)^{1-\theta}$ is the cost function-based productivity. [^2] Otherwise, $B$ is referred to as the dual productivity. See Appendix 1 for more details.
@@
The breakdown of output into investment, capital depreciation, and final consumption is described below:
@@
Here, $\delta\leq 1$ denotes the capital depreciation factor, and $C_{t}$ denotes aggregate consumption in period $t$. Combining equations (1) and (2) and dividing both sides by $L_{t}=L$, which we assume to be constant over time, leads to the following intertemporal dynamics (or roundaboutness) of capital intensity:
@@
where $k_{t}=K_{t}/L$ and $k_{t+1}=K_{t+1}/L$ denote capital intensities in periods $t$ and $t+1$, respectively. Additionally, $y_{t}=Y_{t}/L$ and $c_{t}=C_{t}/L$ denote per capita output and consumption, respectively, in period $t$.
@@ 2.2 Finite horizon dynamic programming
The primary problem here is to solve for an optimal consumption path given a planning horizon. The Bellman equation of the problem is as follows:444 Note that the objective function of the primary problem (2.1) is given at $t=0$, i.e., $\mathcal{V}[n]=\mathcal{V}_{0}[k_{0};n]$.
@@
The optimal trajectory of the state variable $k_{t}$ is given by Lemma 2, which we append to Appendix 2 with proofs, as follows:
@@
The following optimal trajectory of consumption is obtained by (5) and (21) Appendix 2, which must be true according to the proof of Lemma 1.
@@
With formula (7), the population value function is given as follows:
@@
We hereafter aim to maximize this function with respect to the planning horizon $n$.
Our approach to analyzing the population value (8) adopts a numerical rather than an analytical framework because the derivative of $\mathcal{V}[n]$ with respect to $n$ does not seem to provide meaningful insights. In the following section, the optimal trajectory of contributions $\log c_{t}^{*}[n]$ and the population value $\mathcal{V}[n]$, for any given planning horizon $n$, becomes manageable under $\theta=1$ (known as the AK setting) and $\beta<1$. We also find that the trajectory of contributions $\log c_{t}^{*}[\infty]$ and the population value $\mathcal{V}[\infty]$ for an infinite planning horizon is evaluable under $\beta\theta<1$. We therefore base our study of ZD ($\beta=1$) on this parameter setting (i.e., $\beta\theta<1$, which indicates that $\theta<1$, consistent with a Cobb–Douglas model). In the following section, we delve into the abovementioned two broad settings of parameters, namely, AK production with future discounting ($\theta=1$ and $\beta<1$), which we term AK settings, and Cobb–Douglas production without future discounting ($\theta<1$ and $\beta=1$), which we term ZD settings.
Table 1 summarizes the parameter settings chosen for the numerical examinations. Note that cases I–IV correspond to AK settings whose solution paths are characterized by the sign of $\log(A\beta)$, whereas cases V–VII correspond to ZD settings whose solution paths are characterized by the sign of $\log B$. Cases VIII and IX correspond to the parameter settings where $\log({A}\beta)=\log B=0$.
@@
case ${A}$ $\beta$ $\theta$ $\log({A}\beta)$ $\log B$ $n^{*}$ $\mathcal{V}[n^{*}]$ $\mathcal{V}[\infty]$
I $1.012$ $0.992$ $1$ $+$ $\infty$ $84.7$
II $1.01$ $0.992$ $1$ $+$ $95$ $60.0$
III $1/\beta$ $0.992$ $1$ 00 $73$ $55.6$
IV $1.005$ $0.992$ $1$- $58$ $51.0$
@@
    For all cases I–IX, $k_{0}=150$. $B=A(1-\theta)^{1-\theta}\theta^{\theta}$ is the cost-function based productivity.
@@
    $\theta\approx 0.955392$ where 1.2($1{1.2}(1-\theta)^{1-\theta}\theta^{\theta}=B=$ 1.
@@
    $\theta\approx 0.998982$ where $(1/0.992)(1-\theta)^{1-\theta}\theta^{\theta}=B=1$.
@@ 3.1 AK setting
The AK model, which was formally developed by Frankel (1962), is one of the simplest fundamental models of endogenous growth. Here, we employ this production model to study the population value function that discounts future generations to determine whether $n\to\infty$ is an optimal policy. The optimal consumption path (7) for the AK setting with future discounting ($\theta=1$, $\beta<1$) may be specified as follows:
@@
For the sake of the analysis, let us take the derivative with respect to ${n}$.
@@
where $\lambda$ and $\gamma$ are given as follows:
@@
To study the derivative sign of (11), we consider the following two functions, whose equal values, i.e., $f[n]=g[n]$, convey the first-order condition of optimality.
@@
Before we proceed, let us examine the possible range of the parameter $A$. The evaluation of the marginal product of capital (MPK) of our production function (1) net of depreciation, which should coincide with the real interest rate $\rho>0$, viz.,
@@
In AK models with complete capital depreciation, $\theta=\delta=1$ leads to ${A}=1+\rho$. Therefore, assuming that ${A}>1$ is relevant in this setting.555 Note also that $A\beta=(\delta+\rho)\beta$ indicates the discrepancy between the gross interest rate and generational preference rate of the population.
Figure 2 (left) depicts the functions $f[n]$ and $g[n]$ under different parameters for cases I–IV. Clearly, $f[n]$ is a linear function whose slope is $\log({A}\beta)$, and $g[n]$ monotonously increases and approaches zero, i.e., $g[\infty]=0$. Here, we fix the discount factor $\beta$ and differentiate the productivity ${A}$ at four different levels. As long as ${A}\leq 1/\beta$, so that $\log({A}\beta)\leq 0$, i.e., the slope is zero or negative, $f[n]$ will intersect with $g[n]$ at a single point, and the population value function $\mathcal{V}[n]$ has a single peak. If ${A}>1/\beta$ so that $f[n]$ has a positive slope, then $f[n]$ and $g[n]$ could either intersect with two points where the population value $\mathcal{V}[n]$ may rise, fall and then rise again, or never intersect so that the population value may rise indefinitely with respect to the planning horizon $n$. Figure 2 (right) depicts the population value functions for cases I–IV.
Figure 2: Left: The solid lines represent $g[n]$, as defined in (12), for cases I–IV with clear correspondences, i.e., the steepest slope corresponds to case I, while case IV corresponds to the scenario with a negative slope. The dashed line represents $f[n]$, which is also defined in (12). Right: The plots show population value functions $\mathcal{V}[n]$ for cases I–IV with obvious correspondences.
To visualize the optimal trajectory of the variables in various situations, we specify them here in the form of functionals. By referencing (9), the optimal trajectory of the undiscounted contribution to the population value becomes linear with respect to $t$, as follows:
@@
In reference to (21) Appendix 2, the optimal trajectory of capital intensity becomes as follows:
@@
Figure 3 displays the population value function $\mathcal{V}[n]$, specified as (10), for cases I, III, and IV on the top. [^6] These figures correspond to those depicted in Figure 2 (right). Note that, as long as $\beta<1$, the population value function will always converge to a finite value, regardless of the parameter settings in addition to $\beta$. As we let $n\to\infty$ in (10), we have the following:
@@
By comparing the population values obtained at the numerical solution of the first-order condition $f[n]=g[n]$, as referenced in (12), with $\mathcal{V}[\infty]$ from (16), we know the optimal planning horizon $n^{*}$, which we display in Table 1 for cases I–IV. The middle row of Figure 3 displays the optimal trajectories of the undiscounted contributions to the population value, as described in (14), for planning horizons $n=200,400,600$ for cases I, III, and IV (from left to right). Similarly, the bottom row of Figure 3 displays the optimal trajectories of capital intensity, based on (15), for planning horizons $n=200,400,600$ for cases I, III, and IV (from left to right).
Figure 3: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for AK setting cases I (left), III (center), and IV (right). The parameters corresponding to each of the cases are given in Table 1.